Jan 28, 2016 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. Update BCAL clusterizer to improve agreement between simulated and generated MC (Mark)
  2. Modify covariance matrix for FCAL and BCAL to correspond to the current resolution of the detectors (Matt).
  3. Change F250 simulation code to match firmware and add flexibility for new algorithms (Mike)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book
  3. BCAL Action Items and Working Tasks

Tentative Agenda

  1. Announcements
  2. Action Items:
  3. Collaboration Meeting February 18-20, 2016.
  4. (Feb 11) FCAL Insert - Liping Gan
  5. Run Preparations
    1. FCAL Update
    2. BCAL Update
  6. Reconstruction and Simulation
  7. Calibration
    1. JLab and UTFSM SiPM measurements (Elton)
  8. Any other business


Attendees:Noemi, Tegan, Ahmed, Andrei (UofR); Elton, Jon, David, Mark (JLab); Curtis, Will (CMU), Christina, George (Athens)

  1. Announcements
    • Accelerator has requested to close the hall (early) on Monday evening. But they are tracking down some jitter in the injector, which might delay early startup.
  2. Action Items:
  3. Collaboration Meeting February 18-20, 2016.
    • Tegan's talk will be moved to Physics. Overview and Will's talk can be reduced to 20 min.
  4. Run Preparations
    1. FCAL Update
      • Various hardware issues are being sorted out for the Fcal
      • About 35 channels were not reporting signal in the data when the LEDs were flashed.
      • The breakdown of 32 problems is roughly as follows: 14 bases get no HV, 4 bad connectors, 4 bad Fadc channels, 3 base/pmt problems, 1 high switching noise, 6 others need diagnosis.
      • Adesh and Jon are working on fixing the issues before the hall closes.
      • Jon: Manuel mentioned that there are 35 missing entries in the Inventory, which is used to get the HV for the bases. This needs to be filled out
      • Jon: It has been quite some time since cable swaps were checked. The procedure is not fully automatic, so it might take some time.
    2. BCAL Update
      • Bcal is at 5 deg. Elton and Mark tried unsuccessfully yesterday to take LED pulser data. DAQ problems are allegedly fixed and will try this afternoon
      • Beni has tried to take more cosmic data to track down some F125 readout issues, but has run into DAQ problems. He will try again tonight. Previous run list was sent to the list server
  5. Reconstruction and Simulation
    1. David: A new begin run record is being created at the beginning of each run containing hardware run configuration information (FADCs and TDCs). This will make the information accessible from the data file, which previously had to be recovered from either the database or the configuration files.
  6. Calibration
    1. George: Values for veff in the ccdb are about 17 cm/ns from commissioning data. In discussions with George Lolos, who was recently in Athens, they reviewed the data. The results from layer 1 at about 16.5 cm appear to be stable and they are evaluating whether this should be used as the source for ccdb entries. No final decision has been made.
    2. Will: pi0 minimization for fall 2016 see Log Entry 3373005
      • There appears to be some energy dependence to the pi0 peak.
      • Width of pi0 peak is smaller than for the spring data. Could be attributed to the lowering of the NSA samples to 26.
    3. Andrei: converging on new gain calibrations using muons for spring 2015 data (NSA+NSB=60 samples).
      • Expect to show updated results next week
    4. Ahmed: Updated his efficiency report on DocDB
    5. Noemi: Continues to work on calibration with TDCs only, referencing times to other detectors instead of the fadc.
    6. JLab and UTFSM SiPM measurements (Elton)
  7. Any other business